# Lessons learned from field demonstrations of model predictive control and reinforcement learning for residential and commercial HVAC: A review

### v1

@@ 2.3 Reviews of RL for buildings
In response to the growing adoption of RL in various fields such as video games, targeted advertising, robotics, and autonomous vehicles, there has been a marked growth in studies investigating the application of RL for buildings, especially since 2017. The earliest review paper on this topic known to the authors is by Vázquez-Canteli and Nagy [59]. They reviewed 105 research articles that study reinforcement learning algorithms in the context of demand response for smart grids. They emphasized the lack of real-world implementations in the literature, as well as the difficulty of comparing results across different studies due to insufficient standardization of methods, simulation tools, and physical dynamics of the system under consideration. Mason et al. [60] performed a comprehensive review of RL applications for automated building energy management systems. Through their review, the authors found that the specific application of the RL algorithm had a significant impact on the overall energy savings. More specifically, RL algorithms provided savings of around 10% for HVAC appliances, around 20% for water heaters, and greater than 20% for more complete whole building energy management systems. The authors reiterated here that the overwhelming majority of studies were conducted only in simulations. Wang et al. [61] performed a more in-depth review of the formulation of an RL controller for buildings. Their review was sectioned into five major topics of RL problem framework: algorithms, states, actions, rewards, and environment. Some of their key findings were that 77% of the papers reviewed used value-based algorithms such as Q-Learning. They also found that 91% of the studies did not consider historical states, although doing so may violate the Markovian assumptions underlying RL theory, leading to a suboptimal solution. Only 11% of the 77 studies reviewed in this paper were conducted in real buildings. Fu et al. [62] classified different families of RL algorithms and analyzed their suitability to solve different problems related to efficient building control. In contrast to traditional reviews that discuss only RL algorithms that have previously been implemented in other studies, Al Sayed et al. [63] aimed to provide a link between theoretical concepts for RL and the existing gap present in studies for RL applications in buildings. The author concluded that novel techniques such as meta-RL are needed for the periodic extensive retraining of RL agents.
@@
Further topical reviews include Kathirgamanathan et al. [50], who broadly reviewed data-driven predictive control methodologies, including RL and MPC. Michailidis et al. [64] focused specifically on model-free control, while Sierla et al. [65] categorized RL studies based on the decision of action modeling and analyzed the impact of specific choices, such as discrete, binary, or continuous action modeling. Similarly, Shaqour et al. [66] examined the subset of studies that utilized deep reinforcement learning. These modeling reviews are akin to the recent review of Michailidis et al. [67] that surveyed all distributed multi-agent-based building control studies. On the topic of occupant comfort, Han et al. [68] and Chatterjee and Khovalyg [69] explored how comfort is treated in RL studies and identified potential improvements. Lastly, the topical summary by Nagy et al. [70] presented in a ten-question format highlights some of the key modern trends in RL research for buildings. Critical highlights were the limited number of experimental studies, as well as the lack of safety and stability guarantees that some model-based control frameworks can provide.
@@ 2.4 Reviews of MPC and RL field demonstrations
Although numerous papers review MPC and RL applications in HVAC, few focus specifically on field demonstrations. In the first partial review of MPC field demonstrations in commercial buildings, Sturzenegger et al. [14] categorized 10 papers based on the controlled portion of the building, actuator type, experiment duration, and underlying thermal model. Their review revealed gaps in implementation and cost considerations in the existing literature. In addition to their review, they conducted a seven-month MPC experiment in a Swiss office building, showing that MPC provides significant energy savings and comfort improvements, but also requires high implementation costs. Blum et al. [15] investigated 14 field demonstration papers on MPC in commercial buildings. They detailed the key aspects of each paper, such as building type, system type, control variables and objectives, MPC approaches, test periods, and results. They highlighted the diversity of MPC implementations and identified challenges such as deployment complexity and a lack of streamlined tools to facilitate implementation. In addition to their review, they implemented and tested an MPC system in a real office building, proving its feasibility and addressing practical challenges. Focusing on residential buildings, Pergantis et al. [16] summarized 12 field demonstrations of HVAC MPC, including control methods, location of the building, equipment, control action(s), test duration, and performance results. They noted gaps in deployment cost estimates and the near-exclusive focus on energy cost reduction. In addition to their review, they performed a field demonstration in an all-electric house in a cold climate and discussed implementation challenges. The most comprehensive review to date was conducted by Khabbazi et al. [17], who examined 56 field tests in commercial buildings reported in 36 papers between 2005 and 2024. The study covered control methods, building types, objectives, test durations, and reported results, providing information on the variety and evolution of MPC and RL for commercial HVAC. While confirming the ability of MPC and RL to reduce costs and emissions by 15-25% compared with conventional controls, they also noted persistent challenges, such as limited deployment data and scalability issues. More recently, Saloux [18] conducted a partial review, focusing on methods to estimate energy savings from field demonstrations. By evaluating approximately 40 papers covering residential and commercial buildings, the study explored how different benchmarking methods influence reported savings. The findings revealed significant variability in savings estimates depending on the benchmarking method used, emphasizing the need for a more consistent approach to assessing the impact of field-demonstrated energy conservation measures.
@@ 3.1 Search strategy
This paper’s literature review began by identifying studies that used MPC or RL in real-world residential or commercial buildings. Peer-reviewed publications, including journal articles and conference proceedings, were found using search engines such as Google Scholar and Scopus, as well as databases from publishers such as Elsevier, IEEE, Taylor & Francis, Springer and Wiley. When a more detailed journal version of a conference paper was available, the conference paper was excluded. The figures and calculations in this paper do not include theses, dissertations, white papers, or technical reports, although some are cited in the text to recognize their contributions to the field.
To manage the large body of literature, specific search strategies were used. Keywords such as “field demonstration,” “experimental study,” “residential building,” “commercial building,” “model predictive control,” “reinforcement learning,” and “HVAC” were frequently used to refine the search and identify relevant papers. These are only a few of the many keywords used during this process. The identified publications were organized into two tables with approximately 70 columns, each representing a unique characteristic of the study. A full, open-access version of this table can be found in [19]. A contains two more concise versions of these tables, summarizing 24 residential (LABEL:tab:res) and 80 commercial (LABEL:tab:com) papers, respectively. These tables include details such as publication year, control method, building and controlled space description, HVAC system description, test duration, objectives, and reported benefits along with the benchmarking approach.
During the review process, we often contacted authors to gather information that we could not find in the reviewed papers. We contacted authors via email, LinkedIn, and in-person meetings at workshops and conferences. Of the 104 field studies reviewed, 42 (some of which had overlapping authors) required outreach to authors for further clarification. Among these 42, the authors of 36 papers responded, for an 86% response rate. Questions about the other six papers remained unresolved due to, for example, changes in author affiliations or lack of contact information. For these studies, we tried to collect missing data by reading related conference papers, technical reports, theses, and dissertations. We also attempted to estimate parameters such as floor areas and locations ourselves using tools such as Google Maps.
@@ 3.2 Savings evaluation
Occasionally, a reviewed paper presented findings from multiple field tests conducted in multiple building types, regions, and climate zones; or using multiple control methods. When reviewing key results, such as reported benefits, this paper considers the total number of field tests rather than the number of papers. The reported benefits analyzed in the literature include energy and cost savings, comfort improvement, greenhouse gas emissions reduction, peak demand reduction, and more. The primary focus of this paper is on the energy and cost savings achieved by using MPC and RL over conventional control methods. As longer-duration tests tend to be more trustworthy than short-duration tests, this paper emphasizes the duration-weighted average percent savings shown in Eq. 1:
@@ 3.3 Statistical methods
This paper uses the Mann-Whitney U test [71] to compare categorical variables in two groups, such as experiments that controlled a whole-building vs. a subset or building spaces, that used measurement-based vs. simulation-based benchmarks, and that used test durations of less than seven days vs. at least seven days. The Mann-Whitney U test is a nonparametric test of the null hypothesis $H_{0}$ that the populations underlying both categories have the same distribution, vs. the alternative hypothesis $H_{1}$ that the populations have different distributions. We chose the Mann-Whitney U test for its ability to handle data that may not follow Gaussian distributions. To ensure meaningful results, statisticians typically recommend a minimum sample size of 15 per category. For this reason, we do not report test results involving categories with fewer than 15 samples. Due to these constraints, some comparisons could not be tested rigorously, emphasizing the need for better data reporting and more comprehensive field studies in future research.
@@ 4.1 Geography
Fig. 3 shows the distribution of papers by location, classified by country and continent, for residential (left) and commercial (right) buildings. The size of each pie chart corresponds to the total number of papers in its category. Each continent is uniquely color-coded. Both pie charts show that the majority of papers originated from North America and Europe, with the United States and Switzerland leading their respective regions, accounting for approximately 35% and 13% of all papers, respectively. Although some continents are represented in only one category, the combined total of residential and commercial papers includes contributions from all continents. Asia, despite being home to over half of the world’s population, accounts for only 12 of the 104 papers (12%). Other regions are even more underrepresented, with only one paper from Africa, one from South America, and two from Oceania.
Figure 3: Geographical distribution of residential (left) and commercial (right) field demonstration papers by country and continent. Colors indicate continents. Chart sizes are proportional to total numbers of papers.
@@ 4.2 Weather
Fig. 4 shows the distributions of residential (left) and commercial (right) field tests by climate zone. ASHRAE Standard 90.1 [76] defines climate zones based on temperature and humidity characteristics, providing a framework for categorizing regions. The distributions in Fig. 4 are based on the number of tests performed, not the number of papers (as in Fig. 3) because some papers reported multiple tests performed in different regions. The size of each chart corresponds to the total number of tests in that group. Colors indicate temperature zones, while hatching patterns indicate humidity levels.
@@ 4.3.2 Electricity rate structure
Electricity rate structures determine how energy prices vary, ranging from a flat rate (meaning a single, time-invariant price) to dynamic rates (meaning prices that vary with time) such as time-of-use (TOU) or day-ahead pricing. Electricity rate structures influence the potential for cost savings by load shifting. The top center plot in Fig. 5 shows the distribution of tests by electricity rate structure. While residential tests are fairly balanced between flat (27 of 52 tests [52%]) and dynamic rate structures (25 tests [48%]), commercial tests strongly favor flat rates, with 121 of 154 tests (79%) using flat rates and 33 (21%) using dynamic rates.
@@ 4.3.4 Thermal mass of building
Thermal mass affects a building’s ability to store thermal energy, impacting temperature stability and load-shifting effectiveness [77]. Early 21st-century studies examined the role of thermal mass, including optimizing active and passive thermal storage under predictive control [78] and evaluating thermal mass control strategies for cost savings and peak demand reduction [79]. High thermal mass buildings smooth out temperature fluctuations and reduce sensitivity to short-term disturbances [80], while predictive control that activates thermal mass can improve energy efficiency and comfort [81]. The lower left plot of Fig. 5 categorizes studies based on thermal mass into low to moderate, moderate to high, or not reported. Due to a lack of standardized reporting, we classify tests based on available descriptions of the building envelope within the paper itself, previous studies on the same building, relevant institutional or project websites, author correspondence, and, in rare cases, reverse image searches of the building. If none of these methods yielded a reasonable estimate, the thermal mass was marked “not reported.” Despite these efforts, almost one-fifth of residential (11 of 52) and commercial (33 of 154) tests lacked adequate thermal mass information. The reported cases for residential buildings were almost evenly distributed, with 18 tests rated moderate to high thermal mass and 23 as low to moderate. Commercial buildings, on the other hand, showed a stronger preference for higher thermal masses, with 88 tests classified as moderate to high and only 33 as low to moderate. This suggests that commercial structures are more often constructed with heavier materials such as brick and cement, whereas lighter wood-frame construction, while present in both sectors, is more common in residential buildings.
@@ 4.3.6 Control architecture
HVAC control architectures can be classified as high-level or low-level. High-level control typically involves supervisory adjustment of setpoints, such as zone air temperatures or supply air temperatures, relying on device-level controllers such as proportional-integral loops to track high-level setpoints. Low-level control, on the other hand, directly actuates components such as dampers, valves, and compressors, augmenting or replacing the manufacturer’s device-level control logic. The lower right plot of Fig. 5 shows the distribution of tests classified by their control architecture. High-level control was used in 39 (75%) of 52 residential tests and 109 (71%) of commercial tests, while low-level control was used in 13 (25%) and 45 (29%), respectively.
Figure 6: Numbers of residential and commercial tests within each category of protocol-defining characteristics.

### v6

@@ 2.3 Reviews of RL for buildings
In response to the growing adoption of RL in various fields like video games, targeted advertising, robotics, and autonomous vehicles, there has been a marked growth in studies investigating its application in buildings, especially since 2017. The earliest known review on this topic is by Vázquez-Canteli and Nagy [57], who reviewed 105 studies on RL algorithms in the context of demand response for smart grids. They emphasized the lack of real-world implementations in the literature and the difficulty of comparing results across different studies due to insufficient standardization of methods, simulation tools, and system dynamics. Mason and Grijalva [58] reviewed RL applications in building energy management systems. They found that the specific RL application significantly impacted overall energy savings, with savings of around 10% for HVAC, 20% for water heaters, and over 20% for whole-building energy management. The authors reiterated that most studies were conducted only in simulations. Wang and Hong [59] performed a more in-depth review of the formulation of an RL controller for buildings, structuring their review into five key aspects of RL: algorithms, states, actions, rewards, and environment. They found that 77% of the reviewed papers used value-based algorithms like Q-Learning, 91% did not consider historical states, potentially violating the Markovian assumptions underlying RL theory and leading to suboptimal solutions. Also, only about 11% of the 77 studies were conducted in real buildings. Fu et al. [60] classified different families of RL algorithms and analyzed their suitability to solve different problems related to efficient building control. A recent introductory review by Nagy et al. [61] summarized some of the key modern trends in RL research for building energy management in a ten-question format, with an objective to guide its practical adoption and highlight potential future research directions. Critical highlights were the limited number of experimental studies, as well as the lack of safety and stability guarantees that some model-based control frameworks can provide. Unlike traditional reviews that only focus on RL algorithms previously implemented in other studies, Al Sayed et al. [62] sought to provide a link between theoretical RL concepts and the existing gap present in studies for their application in buildings. They concluded that novel techniques such as meta-RL are needed for the periodic extensive retraining of RL agents.
@@
In addition to introductory and comprehensive reviews, several authors published topical reviews and those that discussed RL within broader frameworks of building controls. Kathirgamanathan et al. [54], broadly reviewed data-driven control methodologies, including RL and MPC. With a particular focus on model-free HVAC control, Michailidis et al. [63] presented recent advancements in RL, ANNs, fuzzy logic control, and their hybrids. Sierla et al. [64] categorized RL studies based on the decision of action modeling and analyzed the impact of specific choices, such as discrete, binary, or continuous action modeling. In a review by Shaqour et al. [65], the authors examined the subset of studies that used deep reinforcement learning for various building types, by also discussing research directions and knowledge gaps. These modeling reviews are akin to the recent review of Michailidis et al. [66] that surveyed all distributed multi-agent-based building control studies. On the topic of occupant comfort, Han et al. [67] reviewed RL applications and highlighted its under-use, in particular for indoor air quality and lightning, and the need for more OCC approaches. On a similar topic, Chatterjee and Khovalyg [68] recently discussed the potential of RL to enable dynamic indoor environments that balance energy efficiency and comfort.
@@ 2.4 Reviews of MPC and RL field demonstrations
While many review papers discuss the potential of using MPC and RL in building HVAC control, few specifically focus on field demonstrations. In a partial review of MPC field demonstrations in commercial buildings, Sturzenegger et al. [16] categorized 10 papers based on the controlled portion of the building, actuator type, experiment duration, and underlying thermal model. Their review revealed gaps in implementation and cost considerations in the existing literature. Alongside their review, the authors performed a seven-month MPC experiment in a Swiss office building and demonstrated that MPC provides significant energy savings and comfort improvements; however, it requires high implementation costs. Blum et al. [17] investigated 17 papers on MPC field demonstration in commercial buildings and provided details on the key aspects of each paper, such as building type, system type, control variables and objectives, MPC approaches, test periods, and results. They also developed and tested an MPC system using a Modelica-based [69] workflow in a real office building, proving its feasibility and addressing practical implementation challenges. In addition, they evaluated the implementation effort in person-days and identified barriers involved with integration, data management, and system commissioning. Narrowing their focus to residential cases, Pergantis et al. [18] focused on 12 relevant papers on field demonstration of MPC, covering their control methods, locations, equipments, control actions, test durations, and performance results. Pergantis et al. discussed key research gaps, such as limited information on deployment costs and a narrow focus on energy cost reduction. They deployed their MPC controller in an occupied, all-electric house in a cold climate and discussed implementation challenges. The most comprehensive review to date is that of Khabbazi et al. [6], with a focus on 56 field tests in commercial buildings documented in 36 papers between 2005 and 2024. This study covered key aspects of papers such as control methods, building types, objectives, test durations, and reported results. Khabbazi et al. also introduced duration-weighted averages to assess reported savings and consequently showed the ability of MPC and RL to save energy and costs compared to conventional controllers, with lower savings demonstrated in longer-duration and whole-building tests. The authors highlighted persistent challenges, such as limited deployment data and scalability issues. In a recent work [19], Saloux conducted a partial review focused mainly on MPC and looked at methods for estimating energy savings from field demonstrations. Saloux analyzed about 40 papers on residential and commercial buildings and explored how different benchmarking methods affect reported savings. According to the findings, there is significant variability in savings estimates depending on the benchmarking method, which emphasizes the need for a more consistent assessment of field-demonstrated energy conservation measures.
@@ 3.1 Search strategy
This paper’s literature review began by identifying studies that used MPC or RL in real-world residential or commercial buildings. Peer-reviewed publications, including journal articles and refereed conference proceedings, were found using search engines such as Google Scholar and Scopus, as well as databases from publishers such as Elsevier, IEEE, Taylor & Francis, Springer and Wiley. When a more detailed journal version of a conference paper was available, the conference version was excluded. The figures and calculations in this paper do not include theses, dissertations, white papers, or technical reports.
To manage the large body of literature, specific search strategies were used. Keywords such as “field demonstration,” “experimental study,” “residential building,” “commercial building,” “model predictive control,” “reinforcement learning,” and “HVAC” were frequently used to refine the search and identify relevant papers. These are only a few of the many keywords used during this process. The identified publications were organized into four tables (residential vs. commercial, papers vs. tests), each with a wide range of study characteristics as separate columns. A full, open-access version of this table can be found in [20]. B contains two more concise versions of these tables, summarizing 24 residential (LABEL:tab:res) and 80 commercial (LABEL:tab:com) papers, respectively. These tables include details such as publication year, control method, building and controlled space description, HVAC system description, test durations, objectives, and reported benefits along with the benchmarking approach.
During the review process, we often contacted authors to gather information that we could not find in the reviewed papers. We contacted authors via email, LinkedIn, and in-person meetings at workshops and conferences. Of the 104 field studies reviewed, 45 required outreach to authors for further clarification. We reached out to 32 unique authors (some who wrote multiple papers), receiving responses from 26, with an 81% response rate. A full list of the responding researchers is provided in the Acknowledgments. Questions about the other six papers remained unresolved due to, for example, changes in author affiliations or lack of contact information. For these studies, we tried to collect missing data by reading related conference papers, technical reports, theses, and dissertations. We also attempted to estimate parameters such as floor areas and locations ourselves using tools such as Google Maps.
@@ 3.2 Savings evaluation
Occasionally, a reviewed paper presented findings from multiple field tests conducted in multiple building types, regions, and climate zones; or using multiple control methods. When reviewing key results, such as reported benefits, this paper considers the total number of field tests rather than the number of papers. The reported performance benefits in the literature include energy and cost savings, comfort improvement, greenhouse gas emissions reduction, peak demand reduction, frequency regulation improvement, and more. The main focus of this paper is on the energy and cost savings achieved by using MPC and RL over conventional control methods. Since longer-duration tests tend to be more reliable than short-duration ones, this paper introduces the duration-weighted average percent savings shown in Eq. 1:
@@ 3.3 Statistical methods
This paper uses the Mann-Whitney U test [70] to compare categorical variables in two groups, such as experiments that controlled a whole-building vs. a subset or building spaces, that used measurement-based vs. simulation-based benchmarks, and that used test durations of less than seven days vs. at least seven days. The Mann-Whitney U test is a nonparametric test of the null hypothesis $H_{0}$ that the populations underlying both categories have the same distribution, vs. the alternative hypothesis $H_{1}$ that the populations have different distributions. We chose the Mann-Whitney U test in order to handle data that may not follow Gaussian distributions. Statisticians typically recommend a minimum sample size of 15 per category to guarantee trustworthy results. Consequently, we do not report test results for categories having fewer than 15 samples. Due to these constraints, some comparisons could not be tested rigorously.
@@ 4.1 Geography
Fig. 3 shows the distribution of papers by location, classified by country and continent, for residential (left) and commercial (right) buildings. The size of each pie chart corresponds to the total number of papers in its category. Each continent is uniquely color-coded. Both pie charts show that the bulk of papers originated from North America and Europe, with the United States and Switzerland leading their respective continents with around 35% and 13% of all papers, respectively. While some continents are classified in a single category, the combined total of residential and commercial papers includes contributions from all continents. Despite having over half of the world’s population, Asia accounts for only 12 of the 104 papers (12%). Other regions have much lower representation, with only one paper from Africa, one from South America, and two from Oceania.
Figure 3: Geographical distribution of residential (left) and commercial (right) field demonstration papers by country and continent. Colors indicate continents. Chart sizes are proportional to total numbers of papers. Slices without numbers correspond to a single paper.
@@ 4.2 Weather
Fig. 4 shows the distributions of residential (left) and commercial (right) field tests by climate zone. ASHRAE Standard 90.1 [75] defines climate zones based on temperature and humidity characteristics, which provides a framework for categorizing regions. The distributions in Fig. 4 refer to the number of tests performed, not the number of papers, as opposed to Fig. 3. This difference is because some papers reported multiple tests carried out in different regions. The size of each chart corresponds to the total number of tests in that group. Colors also indicate temperature zones and hatching patterns the humidity levels.
@@ 4.3.2 Electricity rate structure
Electricity rate structures determine how energy prices vary, ranging from a flat rate (meaning a single, time-invariant price) to dynamic rates (meaning prices that vary with time) such as time-of-use or day-ahead pricing. Electricity rate structures influence the potential for cost savings by load shifting. The top center plot in Fig. 5 shows the distribution of tests by electricity rate structure. Residential tests are almost equally balanced between flat (27 of 52 tests, or 52%) and dynamic rate structures (25 of 52 tests, or 48%). In contrast, commercial tests strongly favor flat rates, with 121 of 154 tests (79%) using flat rates and 33 (21%) using dynamic rates.
@@ 4.3.4 Thermal mass of building
Building thermal mass affects temperature stability and load-shifting effectiveness by influencing the building’s ability to store thermal energy [76]. Early studies in the twenty-first century explored the role of thermal mass, such as optimizing active and passive thermal storage under predictive control [77] and evaluating thermal mass control strategies for cost savings and peak demand reduction [78]. Buildings with high thermal mass smooth out temperature fluctuations and reduce sensitivity to short-term disturbances [79]. Additionally, predictive control that activates thermal mass can improve energy efficiency and comfort [80]. The lower left plot of Fig. 5 categorizes studies based on thermal mass into low to moderate, moderate to high, or not reported. Due to a lack of standardized reporting, we classify tests based on available descriptions of the building envelope within the paper itself, previous studies on the same building, relevant institutional or project websites, author correspondence, and, in rare cases, reverse image searches of the building. If none of these methods yielded a reasonable estimate, the thermal mass was marked “not reported.” Despite these efforts, almost one-fifth of residential (11 of 52) and commercial (33 of 154) tests lacked adequate thermal mass information. The reported cases for residential buildings were almost evenly distributed, with 18 tests rated moderate to high thermal mass and 23 as low to moderate. Commercial buildings, however, showed a stronger preference for higher thermal masses, with 88 tests classified as moderate to high and only 33 as low to moderate. This suggests that commercial structures are more often constructed with heavier materials such as brick and cement, whereas lighter wood-frame construction, while present in both sectors, is more common in residential buildings.
@@ 4.3.6 Control architecture
HVAC control architectures can be classified as either high-level or low-level. High-level control typically entails supervisory adjustment of setpoints, such as zone air temperatures or supply air temperatures, with device-level controllers such as proportional-integral loops to track high-level setpoints. Low-level control, however, directly actuates components such as dampers, valves, and compressors, augmenting or replacing the manufacturer’s device-level control logic. The lower right plot of Fig. 5 shows the distribution of tests classified by their control architecture. High-level control was used in 39 (75%) of 52 residential tests and 109 (71%) of commercial tests, while low-level control was used in 13 (25%) and 45 (29%), respectively.
4.4 Algorithm implementation choices
This section summarizes the specific algorithm implementation choices that most researchers made in the field papers reviewed here. For tutorial reviews of the MPC and RL algorithms and implementation options in the context of HVAC, we refer the reader to [42, 43, 44] for MPC and [57, 59] for RL.
4.4.1 MPC implementation
The basic components of an MPC algorithm are the system model; a method for forecasting uncertain model inputs; an optimization model including decision variables, an objective function, and constraint functions; and a numerical optimization solver. When state variables are imperfectly observed, MPC is typically combined with a recursive state estimator, such as a Kalman filter for linear systems, or for nonlinear systems, a particle filter or an extended or unscented Kalman filter.
Two-thirds of the MPC field papers used linear models of the building dynamics. Most models followed a thermal circuit structure (also known as an RC network), with a typical model order of one to three states per thermal zone. Linear autoregressive time-series models were also fairly common. Smaller numbers of papers modeled system dynamics using high-dimensional, nonlinear building simulation software such as EnergyPlus [82], TRNSYS [83], or Modelica. Other papers used nonlinear regression models, such as neural networks or regression trees. Most papers used time step durations of five to 60 minutes. HVAC equipment performance curves were typically modeled using polynomials of degree one to three, often fit to manufacturer performance data, although a few papers used lookup tables or neural networks.
The MPC field papers used a wide range of methods for forecasting uncertain model inputs such as internal heat gains from sunlight, electrical loads, and bodies. Popular forecasting methods include linear autoregressive time-series models and neural networks. In addition to implementing their own forecasting methods, many papers also downloaded external forecasts of weather variables or energy prices. Most papers used forecast horizons of two to 24 hours.
The MPC field papers often minimized the cumulative input energy to HVAC equipment over the forecast horizon or the cumulative energy cost. A small number of papers aimed, additionally or instead, to minimize the peak electricity demand over the forecast horizon, or to maximize revenues from providing demand response or power grid reliability services. Many papers augmented the objective function with ‘soft constraints,’ meaning constraint violation penalties. Indoor air temperature constraints were often ‘softened’ to avoid infeasible optimization problem instances. Most papers enforced HVAC equipment limits via (hard) constraints, rather than (soft) penalties. Indoor temperatures and heat transfer rates were common optimization variables, although some papers directly optimized equipment set-points, such as air or water flow rates, supply temperatures, compressor speeds, or fan speeds.
After combining dynamical models, forecasts, and optimization models, the MPC field papers arrived at a wide range of optimization problem structures. Many papers solved convex optimization problems – typically linear or quadratic programs – for which off-the-shelf solvers can return globally optimal solutions in polynomial time [84]. Many other papers solved nonconvex optimization problems, with nonconvexities typically arising from nonlinear dynamics, nonconvex equipment performance curves, or integer-valued decision variables representing, for example, equipment on/off states. Most nonconvex optimization solvers cannot guarantee global solution in polynomial time, but often return locally optimal solutions that significantly improve upon user-specified initial guesses.
4.4.2 RL implementation
The basic components of an RL algorithm are the state and action spaces, reward function, and exploration method. In HVAC applications, deploying an untrained RL agent in hardware could put equipment at risk or require long training times before convergence to an effective policy. For these reasons, all RL field papers reviewed here pre-trained an agent offline in a simulator before deploying it in hardware. Implemented in this way, even model-free RL variants require a system model for offline pre-training. The most common system models used in RL field papers were neural networks, followed by EnergyPlus and the low-order linear thermal circuits that are popular in MPC.
Nearly all RL field papers used the Q-learning or actor-critic RL variants. Typical actions included equipment on/off states, indoor air temperature set-points, supply fluid flow or temperature set-points, or fan speeds. Nearly all RL papers included indoor air temperatures as states; many also included states related to the time of day, weather measurements or forecasts, and occupancy measurements.
Like the objective function in MPC, the RL reward function encodes user-specified goals. As with MPC, most RL field papers aimed to reduce input energy or energy costs. Most RL variants do not readily accommodate constraints, so papers typically augmented the reward function with violation penalties associated with thermal comfort or equipment limits. Several papers also rewarded temporal smoothness of the action signal.
Exploration methods balance exploration/exploitation trade-offs by steering the system toward state-action configurations that may be under-explored. The most popular exploration methods in HVAC RL field papers were the $\epsilon$-greedy and Boltzmann approaches. A few papers encouraged exploration by adding Gaussian noise to the implemented actions.